OpenStudy (anonymous):

Divide and Simplify : x^2-x-6/x^2-2x-3 divide by x^2+x-12/x^2+5x-12 PLEASE HELPPPPP !! I have to turn this in ASAP Show work please

OpenStudy (anonymous):

when dividing by a fraction, it is the same as multiplying by its reciprocal \[\frac{x}{y} \div \frac{m}{n} = \frac{x}{y} * \frac{n}{m}\]